2x-4y=8 What is the slope interception?
Answer by checkley79(3341)

Y=mX+b where m=the slope & b=the Y intercept.
2x-4y=8
-4y=-2x+8
y=-2x/-4+8/-4
y=x/2-2 This line has a slope=(1/2) & a Y intercept=(-2).